Parth (parthkohli):

Maximum and minimum values of the expression \(y = \cos \theta \left(\sin \theta + \sqrt{\sin^2 \theta + \sin^2 \alpha}\right)\)?

Parth (parthkohli):

@ganeshie8 So I considered the general rule of thumb, i.e., converting the expression to \(a \cos t + b\sin t\). Identities that come to my mind are \(\sin^2 \theta = \frac{1 - \cos 2\theta}{2}\) and \(\cos \theta \sin \theta = \frac{1}{2} \sin 2\theta\).

mathslover (mathslover):

Well, did you get any noticeable thing by applying the above identities?

Parth (parthkohli):

I didn't...

Parth (parthkohli):

But I'm pretty sure that I should.

mathslover (mathslover):

Cool. I'll give it a try then.

Parth (parthkohli):

My main concern is removing the stupid sqrt.

Parth (parthkohli):

Or do I have to do that? Maybe I can manage with the sqrt.

mathslover (mathslover):

I'm getting something similar to this : \(\cfrac{1}{2} \sin 2 \theta + \cos \sqrt{1 - \cos (\theta + \alpha) \cos (\theta - \alpha) } \) (After simplification)

Parth (parthkohli):

same here

mathslover (mathslover):

Hmm! What if we square both sides? Will that help us?

Parth (parthkohli):

that may be a good idea

mathslover (mathslover):

Hmm! It is getting more complicated if we square both sides, I guess. :/

mathslover (mathslover):

Should we attempt for a Calculus method?

Parth (parthkohli):

sure.

ganeshie8 (ganeshie8):

*

mathslover (mathslover):

@ganeshie8 - Help please :D

Parth (parthkohli):

looks like I am getting to something through the squaring thing

mathslover (mathslover):

Cool!

Parth (parthkohli):

Yup, getting a quadratic expression in \(\cos 2\theta.\)

mathslover (mathslover):

Oh, great. Good work

Parth (parthkohli):

\[y^2 = \dfrac{-2\cos^2 2\theta + 2\sin^2 \alpha \cos 2 \theta +( 2\sin^2\alpha + 1 )}{4}\]

Parth (parthkohli):

The problem with squaring is that (a) it took me a while to find this expression and (b) even though squaring can help me with the max, it can't help me with the min.

Parth (parthkohli):

wait is that expression even right? >_<

Parth (parthkohli):

OK, simple fix: add 1 to the numerator.

Parth (parthkohli):

Whoops, I meant subtract. Yes. Too many mistakes... =(

Parth (parthkohli):

\[y^2 = \dfrac{-2 \cos^2 2\theta + 2\sin^2\alpha \cos 2\theta + 2\sin^2 \alpha }{4}\]

Parth (parthkohli):

Scratch that. I meant add.

Parth (parthkohli):

what is wrong with me

Parth (parthkohli):

Can someone please suggest an easier way to do this? :(

OpenStudy (anonymous):

used vectors..